# A Hausdorff-measure boundary element method for acoustic scattering by fractal screens

### v1

@@ 1 Introduction
A classical problem in the study of acoustic, electromagnetic and elastic wave propagation is the scattering of a time-harmonic incident wave by an infinitesimally thin screen (or “crack”). In the simplest configuration the incident wave propagates in $\mathbb{R}^{n+1}$ (typically $n=1,2$) and the screen $\Gamma$ is assumed to be a bounded subset of the hyperplane $\Gamma_{\infty}=\mathbb{R}^{n}\times\{0\}$. In standard analyses the set $\Gamma$ is assumed (either explicitly or implicitly) to be a relatively open subset of $\Gamma_{\infty}$ with smooth relative boundary $\partial\Gamma$. But in a recent series of papers [14, 12, 16, 2] it has been shown how well-posed boundary value problems (BVPs) and associated boundary integral equations (BIEs) for the acoustic version of this screen problem (with either Dirichlet, Neumann or impedance boundary conditions) can be formulated, analysed and discretized for arbitrary screens with no regularity assumption on $\Gamma$. In particular, this encompasses situations where either $\partial\Gamma$ or $\Gamma$ itself has a fractal nature. The study of wave scattering by such fractal structures is not only interesting from a mathematical point of view, but is also relevant for numerous applications including the scattering of electromagnetic waves by complex ice crystal aggregates in weather and climate science [39] and the modelling of fractal antennas in electrical engineering [45].
@@
Our focus in this paper is on the Dirichlet (sound soft) acoustic scattering problem in the case where $\Gamma$ itself is fractal. [^1] We note that our methods and results apply, with obvious modifications, to the analogous (yet simpler) problem in potential theory, in which the Helmholtz equation is replaced by the Laplace equation. We shall assume throughout that $\Gamma$ is a compact $d$-set (see §2.1 for the definition) for some $n-1<d\leq n$, which in particular implies that $\Gamma$ has Hausdorff dimension equal to $d$. More specifically, our attention will be on the special case where $\Gamma$ is the self-similar attractor of an iterated function system of contracting similarities, in particular on the case where $\Gamma$ satisfies a certain disjointness condition (described in §2.3), in which case $\Gamma$ has (as a subset of $\mathbb{R}^{n}$) empty interior and zero Lebesgue measure. An example in the case $n=1$ is the middle-third Cantor set, which is a $d$-set for $d=\log{2}/\log{3}$; an example in the case $n=2$ is the middle-third Cantor dust shown in Figure 1, which is a $d$-set for $d=\log{4}/\log{3}$. For such $\Gamma$, well-posed BVP and BIE formulations for the Dirichlet scattering problem were analysed in [12], where it was shown that the exact solution of the BIE lies in the function space $H^{-1/2}_{\Gamma}=\{u\in H^{-1/2}(\Gamma_{\infty}):\operatorname{supp}{u}\subset\Gamma\}$ [12, §3.3]. The assumption that d>n1d>n-1 implies that this space is non-trivial, and that for non-zero incident data the BIE solution is non-zero, so the screen produces a non-zero scattered field. Our aim in this paper is to develop and analyse a boundary element method (BEM) that can efficiently compute this BIE solution.
@@
One obvious approach, adopted in [16] (and see also [29, 34, 2]), is to apply a conventional BEM on a sequence of smoother (e.g. Lipschitz) “prefractal” approximations to the underlying fractal screen, such as those illustrated in Figure 1 for the middle-third Cantor dust. (In this example each prefractal is a union of squares.) When $\Gamma$ has empty interior (as in the current paper) this is necessarily a “non-conforming” approach, in the sense that the resulting discrete approximations do not lie in $H^{-1/2}_{\Gamma}$, the space in which the continuous variational problem is posed. This is because conventional BEM basis functions are elements of $L_{2}(\Gamma_{\infty})$, the intersection of which with $H^{-1/2}_{\Gamma}$ is trivial. This complicates the analysis of Galerkin implementations, since Céa’s lemma (e.g., [40, Theorem 8.1]), or its standard modifications, cannot be invoked. In [16] we showed how this can be overcome using the framework of Mosco convergence, proving that, in the case of piecewise-constant basis functions, the BEM approximations on the prefractals converge to the exact BIE solution on $\Gamma$ as the prefractal level tends to infinity, provided that the prefractals satisfy a certain geometric constraint and the corresponding mesh widths tend to zero at an appropriate rate [16, Thm. 5.3]. However, while [16] provides, to the best of our knowledge, the first proof of convergence for a numerical method for scattering by fractals, we were unable in [16] to prove any rates of convergence.
@@
Finally, in §6, we present numerical results, for cases where $\Gamma$ is a Cantor set or Cantor dust, illustrating the sharpness of our theoretical predictions. We show that our error estimates appear to apply also in cases, such as the Sierpinski triangle, where $\Gamma$ is not disjoint so that the conditions of our theory are not fully satisfied. We also make comparisons, in terms of accuracy as a function of numbers of degrees of freedom, with numerical results obtained by applying conventional BEM on a sequence of prefractal approximations to $\Gamma$, for which we have, as discussed above, only a much more limited theory [16].
@@ 2.1 Hausdorff measure and dimension
As in [31, §1.1] and [43, §3], given $0<d\leq n$, a closed set $\Gamma\subset\mathbb{R}^{n}$ is said to be a $d$-set if there exist $c_{2}>c_{1}>0$ such that
@@
where $B_{r}(x)\subset\mathbb{R}^{n}$ denotes the closed ball of radius $r$ centred on $x$. Condition (1) implies that $\Gamma$ is uniformly locally $d$-dimensional in the sense that ${\rm dim_{H}}(\Gamma\cap B_{r}(x))=d$ for every $x\in\Gamma$ and $r>0$. In particular (see the discussion in [24, §2.4]) (1) implies that $0<\mathcal{H}^{d}(\Gamma\cap B_{R}(0))<\infty$ for all sufficiently large $R>0$, so that ${\rm dim_{H}}(\Gamma)=d$.
In the case $d=n$ an open set can satisfy (1). We call $\Omega\subset\mathbb{R}^{n}$ an open $n$-set if $\Omega$ is open and there exists $c_{1}>0$ such that
$\left|\Omega\cap B_{r}(x)\right|\geq c_{1}r^{n},\qquad x\in\Omega,\quad 0<r\leq 1.$ (2)
If $\Omega$ is an open $n$-set then (see [31, Prop. 1 on p. 205]) $\Gamma:=\overline{\Omega}$ is an $n$-set and $|\partial\Omega|=0$, so that also $|\partial\Gamma|=0$. Conversely, if $\Gamma$ is an $n$-set and $|\partial\Gamma|=0$, it is easy to see that $\Gamma=\overline{\Omega}$, where $\Omega:=\Gamma^{\circ}$, the interior of $\Gamma$, and that $\Omega$ is an open $n$-set. We note that open $n$-sets are also known as interior regular domains, see [8, Def. 5.7].
@@ 2.2 Singular integrals on compact $d$-sets
Our Hausdorff BEM involves the discretization of a weakly singular integral equation in which integration is carried out with respect to Hausdorff measure. In order to derive the basic integrability results we require, we appeal to the following lemma, which is [9, Lemma 2.13] with the dependence of the equivalence constants made explicit.
@@
If $\Gamma\subset\mathbb{R}^{n}$ is compact and the right-hand inequality in (1) holds, i.e., $\mathcal{H}^{d}(\Gamma\cap B_{r}(x))\leq c_{2}r^{d}$, for $x\in\Gamma$, $0<r\leq 1$, then, following the proof of [9, Lemma 2.13], we see that the right hand bound in (3) holds, with $C_{2}$ depending only on $c_{2}$, $n$, and the diameter of $\Gamma$.
@@
Corollary 2.3(i) is related to the more general correspondence between Hausdorff dimension and so-called “capacitary dimension” - see, e.g., [43, §17.11].
@@ 2.3 Iterated function systems
We shall assume throughout that $\Gamma$ satisfies the open set condition (OSC) [24, (9.11)], meaning that there exists a non-empty bounded open set $O\subset\mathbb{R}^{n}$ such that
@@
Then [43, Thm. 4.7] $\Gamma$ is a $d$-set, where $d\in(0,n]$ is the unique solution of
@@
Returning to the general, not necessarily homogeneous, case, the OSC (5) also implies (again, see [43, Thm. 4.7]) that $\Gamma$ is self-similar in the sense that the sets
@@
which are congruent copies of $\Gamma$, satisfy
@@
That is, $\Gamma$ can be decomposed into $M$ congruent copies of itself, whose pairwise intersections have Hausdorff measure zero. For many of our results we make the additional assumption that the sets $\Gamma_{1},\ldots,\Gamma_{M}$ are disjoint, in which case we say that the IFS attractor $\Gamma$ is disjoint. We recall that if $\Gamma$ is disjoint then it is totally disconnected [24, Thm. 9.7]. Examples of IFS attractors that are disjoint are the Cantor set and Cantor dust (Figure 1), while an example that satisfies the OSC but is not disjoint is the Sierpinski triangle. The next result relates disjointness to the OSC.
@@
Arguing as on [16, p. 809], $\Gamma\subset\overline{O}$, and, if $d<n$, then the Lebesgue measure of $s(\overline{O})$, $|s(\overline{O})|\leq\sum_{m=1}^{M}\rho_{m}^{n}|\overline{O}|<\sum_{m=1}^{M}\rho_{m}^{d}|\overline{O}|=|\overline{O}|$, so that $s(\overline{O})\neq\overline{O}$, so that (since $\Gamma=s(\Gamma)$), $\Gamma\neq\overline{O}$.
Suppose now that $d=n$. Arguing as above and on [16, p. 809], $|s(O)|=\sum_{m=1}^{M}|s_{m}(O)|=\sum_{m=1}^{M}\rho_{m}^{n}|O|=|O|$, so that $|s(O)|=|O|$. As claimed in [16, p. 809], this implies that $s(\overline{O})=\overline{O}$. To see this, note that $s(O)\subset O$, so that $s(\overline{O})=\overline{s(O)}\subset\overline{O}$. Thus, if $s(\overline{O})\neq\overline{O}$, there exists $x\in\overline{O}\setminus s(\overline{O})$ so that, for some $y\in O$ near $x$ and some $\epsilon>0$, $B_{\epsilon}(y)\subset O\setminus s(\overline{O})\subset O\setminus s(O)$, which contradicts $|O|=|s(O)|$. Further, since $\Gamma$ is the unique fixed point of $s$, $s(\overline{O})=\overline{O}$ implies $\Gamma=\overline{O}$, and that $\Gamma$ is not disjoint then follows from Lemma 2.5. ∎
@@ 2.4 Function spaces on subsets of $\mathbb{R}^{n}$
Here we collect some results on function spaces from [14, 8, 30]. Our function spaces will be complex-valued, and we shall repeatedly use the following terminology relating to dual spaces [^3] For notational convenience we follow the convention of [14, 16] and work throughout with dual spaces of antilinear functionals rather than linear functionals.. If $X$ and $Y$ are Hilbert spaces, $X^{*}$ is the dual space of $X$, and $I:Y\to X^{*}$ is a unitary isomorphism, we say that $(Y,I)$ is a unitary realisation of $X^{*}$ and define the duality pairing $\langle y,x\rangle_{Y\times X}:=Iy(x)$, for $y\in Y,x\in X$. Having selected a unitary realisation $(Y,I)$ of $X^{*}$, we adopt $(X,I^{*})$ as our unitary realisation of $Y^{*}$, where $I^{*}x(y):=\overline{\langle y,x\rangle_{Y\times X}}$, so that $\langle x,y\rangle_{X\times Y}=\overline{\langle y,x\rangle_{Y\times X}}$, for $y\in Y,x\in X$.
@@
In practice we shall view $\mathbb{L}_{2}(\Gamma)$ as a space of functions on $\Gamma$, by identifying elements of $\mathbb{L}_{2}(\Gamma)$ with their restrictions to $\Gamma$. The dual space $(\mathbb{L}_{2}(\Gamma))^{*}$ can be realised in the standard way as $(\mathbb{L}_{2}(\Gamma),\mathbb{I})$, where $\mathbb{I}:\mathbb{L}_{2}(\Gamma)\to\mathbb{L}_{2}(\Gamma)^{*}$ is the Riesz map (a unitary isomorphism) defined by $\mathbb{I}f(\tilde{f})=(f,\tilde{f})_{\mathbb{L}_{2}(\Gamma)}=\int_{\Gamma}f(x)\overline{\tilde{f}(x)}\,\mathrm{d}\mathcal{H}^{d}(x)$.
Now assume that $\Gamma$ is a $d$-set and that $0<d\leq n$. Then function spaces on $\mathbb{R}^{n}$ and $\Gamma$ are related via the trace operator $\mathrm{tr}_{\Gamma}$ of [43, §18.5]. Defining $\mathrm{tr}_{\Gamma}(\varphi)=\varphi|_{\Gamma}\in\mathbb{L}_{2}(\Gamma)$ for $\varphi\in C_{0}^{\infty}(\mathbb{R}^{n})$, one can show [43, Thm 18.6] that if
@@
with dense range. [^5] For $0<d<n$ the operator $\mathrm{tr}_{\Gamma}$ extends further to a surjective map $\mathrm{tr}_{\Gamma}:B^{\frac{n-d}{2}}_{2,1}(\mathbb{R}^{n})\to\mathbb{L}_{2}(\Gamma)$ on the Besov space $B^{\frac{n-d}{2}}_{2,1}(\mathbb{R}^{n})$ (in which $H^{s}{(\mathbb{R}^{n})}$ is dense), but we do not need that generality here. For details see [8, §6] and [43, Thm 18.6]. This trivially holds also for $d=n$, since the embedding of $H^{s}(\mathbb{R}^{n})$ into $L_{2}(\mathbb{R}^{n})$, for $s>0$, and the trace $\mathrm{tr}_{\Gamma}:L_{2}(\mathbb{R}^{n})\to\mathbb{L}_{2}(\Gamma)$ are both continuous with dense range. Setting
@@
for $t^{\prime}>t>0$, and the embeddings are continuous with dense range. As explained in [8, Rem. 6.4] ($\mathbb{H}^{t}(\Gamma)$ is denoted $\mathbb{H}^{t}_{2,0}(\Gamma)$ in [8, §6]), for the case $0<d<n$, under the further assumption that $t<1$, $\mathbb{H}^{t}(\Gamma)$ coincides (with equivalent norms) with the Besov space $B^{t}_{2,2}(\Gamma)$ of [31]. Arguing in the same way, using [31, Theorem VI.1] and that $H^{s}(\mathbb{R}^{n})$ coincides with the Besov space $B^{s}_{2,2}(\mathbb{R}^{n})$ (e.g., [31, p. 8]) [^6] Our standard notation for Besov spaces on $\mathbb{R}^{n}$ is that of, e.g., [44]., this holds also for $d=n$.
In the particular case that $\Gamma=\overline{\Omega}$ and $\Omega$ is an open $n$-set in $\mathbb{R}^{n}$, it holds that [^7] To see this, note that $\mathbb{H}^{s}(\Gamma)$ and $H^{s}(\Gamma)$ are unitarily isomorphic to the quotient spaces $H^{s}(\mathbb{R}^{n})/\ker(\mathrm{tr}_{\Gamma})$ and $H^{s}(\mathbb{R}^{n})/\ker(|_{\Omega})$, respectively, where $\mathrm{tr}_{\Gamma}:H^{s}(\mathbb{R}^{n})\to\mathbb{L}_{2}(\Gamma)$ and $|_{\Omega}:H^{s}(\mathbb{R}^{n})\to L_{2}(\Omega)$, and that, if $|\partial\Omega|=0$ (which holds if $\Omega$ is an open $n$-set), then these quotient spaces coincide (with equality of norms) since $\ker(\mathrm{tr}_{\Gamma})=\ker(|_{\Omega})$. This last claim follows from the fact that the operation of restriction to $\Omega$ is an isomorphism from $\mathbb{L}_{2}(\Gamma)$ onto $L_{2}(\Omega)$ if $|\partial\Omega|=0$ (recall that $\mathcal{H}^{n}$ is a constant times the Lebesgue measure).
$\mathbb{H}^{s}(\Gamma)=H^{s}(\Omega),\qquad s>0,$ (12)
in the sense that the operation of restriction to $\Omega$ establishes a one-to-one correspondence from $\mathbb{H}^{s}(\Gamma)$ onto $H^{s}(\Omega)$, the norms of corresponding elements being the same.
@@
2.5 Function spaces on screens

### v2

@@ 1 Introduction
A classical problem in the study of acoustic, electromagnetic and elastic wave propagation is the scattering of a time-harmonic incident wave by an infinitesimally thin screen (or “crack”). In the simplest configuration the incident wave propagates in $\mathbb{R}^{n+1}$ (typically $n=1,2$) and the screen $\Gamma$ is assumed to be a bounded subset of the hyperplane $\Gamma_{\infty}=\mathbb{R}^{n}\times\{0\}$. In standard analyses the set $\Gamma$ is assumed (either explicitly or implicitly) to be a relatively open subset of $\Gamma_{\infty}$ with smooth relative boundary $\partial\Gamma$. But in a recent series of papers [17, 15, 19, 3] it has been shown how well-posed boundary value problems (BVPs) and associated boundary integral equations (BIEs) for the acoustic version of this screen problem (with either Dirichlet, Neumann or impedance boundary conditions) can be formulated, analysed and discretized for arbitrary screens with no regularity assumption on $\Gamma$. In particular, this encompasses situations where either $\partial\Gamma$ or $\Gamma$ itself has a fractal nature. The study of wave scattering by such fractal structures is not only interesting from a mathematical point of view, but is also relevant for numerous applications including the scattering of electromagnetic waves by complex ice crystal aggregates in weather and climate science [46] and the modelling of fractal antennas in electrical engineering [52]. In applications the physical object generally only exhibits a certain number of levels of fractal structure; nonetheless, fractals provide an idealised mathematical model for objects that have self-similar structure at multiple lengthscales.
@@
Our focus in this paper is on the Dirichlet (sound soft) acoustic scattering problem in the case where $\Gamma$ itself is fractal. [^1] We note that our methods and results apply, with obvious modifications, to the analogous (yet simpler) problem in potential theory, in which the Helmholtz equation is replaced by the Laplace equation. We shall assume throughout that, for some $n-1<d\leq n$, $\Gamma$ is a compact $d$-set (i.e., $\Gamma$ is compact as a subset of $\Gamma_{\infty}$ and is a $d$-set as defined in §2.1) which in particular implies that $\Gamma$ has Hausdorff dimension equal to $d$. More specifically, our attention will be on the special case where $\Gamma$ is the self-similar attractor of an iterated function system of contracting similarities, in particular on the case where $\Gamma$ satisfies a certain disjointness condition (described in §2.3), in which case $\Gamma$ has (as a subset of $\mathbb{R}^{n}$) empty interior and zero Lebesgue measure. An example in the case $n=1$ is the middle-third Cantor set, which is a $d$-set for $d=\log{2}/\log{3}$; an example in the case $n=2$ is the middle-third Cantor dust shown in Figure 1, which is a $d$-set for $d=\log{4}/\log{3}$. For such $\Gamma$, well-posed BVP and BIE formulations for the Dirichlet scattering problem were analysed in [15], where it was shown that the exact solution of the BIE lies in the function space $H^{-1/2}_{\Gamma}=\{u\in H^{-1/2}(\Gamma_{\infty}):\operatorname{supp}{u}\subset\Gamma\}$ [15, §3.3]. The assumption that d>n1d>n-1 implies that this space is non-trivial, and that for non-zero incident data the BIE solution is non-zero, so the screen produces a non-zero scattered field. Our aim in this paper is to develop and analyse a boundary element method (BEM) that can efficiently compute this BIE solution.
@@
One obvious approach, adopted in [19] (and see also [35, 40, 3]), is to apply a conventional BEM on a sequence of smoother (e.g. Lipschitz) “prefractal” approximations to the underlying fractal screen, such as those illustrated in Figure 1 for the middle-third Cantor dust [^2] For recent overviews of the conventional BEM literature for Lipschitz or smoother screens see [19] or [2, 21, 34].. (In this example each prefractal is a union of squares.) When $\Gamma$ has empty interior (as in the current paper) this is necessarily a “non-conforming” approach, in the sense that the resulting discrete approximations do not lie in $H^{-1/2}_{\Gamma}$, the space in which the continuous variational problem is posed. This is because conventional BEM basis functions are elements of $L_{2}(\Gamma_{\infty})$, the intersection of which with $H^{-1/2}_{\Gamma}$ is trivial. This complicates the analysis of Galerkin implementations, since Céa’s lemma (e.g., [47, Theorem 8.1]), or its standard modifications, cannot be invoked. In [19] we showed how this can be overcome using the framework of Mosco convergence, proving that, in the case of piecewise-constant basis functions, the BEM approximations on the prefractals converge to the exact BIE solution on $\Gamma$ as the prefractal level tends to infinity, provided that the prefractals satisfy a certain geometric constraint and the corresponding mesh widths tend to zero at an appropriate rate [19, Thm. 5.3]. However, while [19] provides, to the best of our knowledge, the first proof of convergence for a numerical method for scattering by fractals, we were unable in [19] to prove any rates of convergence.
@@
In §6 we present numerical results, for cases where $\Gamma$ is a Cantor set or Cantor dust, illustrating the sharpness of our theoretical predictions. We show that our error estimates appear to apply also in cases, such as the Sierpinski triangle, where $\Gamma$ is not disjoint so that the conditions of our theory are not fully satisfied. We also make comparisons, in terms of accuracy as a function of numbers of degrees of freedom, with numerical results obtained by applying conventional BEM on a sequence of prefractal approximations to $\Gamma$, for which we have, as discussed above, only a much more limited theory [19].
In §7 we offer some conclusions and suggestions for future work. In Appendix C we provide a table of definitions for easy reference.
@@ 2.1 Hausdorff measure and dimension
As in [37, §1.1] and [50, §3], given $0<d\leq n$, a closed set $\Gamma\subset\mathbb{R}^{n}$ is said to be a $d$-set if there exist $c_{2}>c_{1}>0$ such that
@@
where $B_{r}(x)\subset\mathbb{R}^{n}$ denotes the closed ball of radius $r$ centred on $x$. Condition (1) implies that $\Gamma$ is uniformly locally $d$-dimensional in the sense that ${\rm dim_{H}}(\Gamma\cap B_{r}(x))=d$ for every $x\in\Gamma$ and $r>0$. In particular (see the discussion in [28, §2.4]) (1) implies that $0<\mathcal{H}^{d}(\Gamma\cap B_{R}(0))<\infty$ for all sufficiently large $R>0$, so that ${\rm dim_{H}}(\Gamma)=d$.
@@ 2.2 Singular integrals on compact $d$-sets
Our Hausdorff BEM involves the discretization of a weakly singular integral equation in which integration is carried out with respect to Hausdorff measure. In order to derive the basic integrability results we require, we appeal to the following lemma, which is [12, Lemma 2.13] with the dependence of the equivalence constants made explicit.
@@
If $\Gamma\subset\mathbb{R}^{n}$ is compact and the right-hand inequality in (1) holds, i.e., $\mathcal{H}^{d}(\Gamma\cap B_{r}(x))\leq c_{2}r^{d}$, for $x\in\Gamma$, $0<r\leq 1$, then, following the proof of [12, Lemma 2.13], we see that the right-hand bound in (2) holds, with $C_{2}$ depending only on $c_{2}$, $n$, and the diameter of $\Gamma$.
@@
Corollary 2.3(i) is related to the more general correspondence between Hausdorff dimension and so-called “capacitary dimension” - see, e.g., [50, §17.11].
@@ 2.3 Iterated function systems
We shall assume throughout that $\Gamma$ satisfies the open set condition (OSC) [28, (9.11)], meaning that there exists a non-empty bounded open set $O\subset\mathbb{R}^{n}$ such that
@@
Then [50, Thm. 4.7] $\Gamma$ is a $d$-set, where $d\in(0,n]$ is the unique solution of
@@
Returning to the general, not necessarily homogeneous, case, the OSC (4) also implies (again, see [50, Thm. 4.7]) that $\Gamma$ is self-similar in the sense that the sets
@@
which are similar copies of $\Gamma$, satisfy
@@
That is, $\Gamma$ can be decomposed into $M$ similar copies of itself, whose pairwise intersections have Hausdorff measure zero. For many of our results we make the additional assumption that the sets $\Gamma_{1},\ldots,\Gamma_{M}$ are disjoint, in which case we say that the IFS attractor $\Gamma$ is disjoint. We recall that if $\Gamma$ is disjoint then it is totally disconnected [28, Thm. 9.7]. Examples of IFS attractors that are disjoint are the Cantor set (§6.1) and Cantor dust (Figure 1 and §6.2), while examples that satisfy the OSC but are not disjoint include the Sierpinski triangle (§6.5(ii)) and the unit interval $[0,1]$. The latter is the attractor of the IFS (124) with $\rho=1/2$, showing that IFS attractors, while self-similar, need not be fractal. A compendium of well-known fractal IFS attractors can be found at [41].
The next result relates disjointness to the OSC.
@@
Arguing as on [19, p. 809], $\Gamma\subset\overline{O}$, and, if $d<n$, then the Lebesgue measure of $s(\overline{O})$, $|s(\overline{O})|\leq\sum_{m=1}^{M}\rho_{m}^{n}|\overline{O}|<\sum_{m=1}^{M}\rho_{m}^{d}|\overline{O}|=|\overline{O}|$, so that $s(\overline{O})\neq\overline{O}$, so that (since $\Gamma=s(\Gamma)$), $\Gamma\neq\overline{O}$.
Suppose now that $d=n$. Arguing as above and on [19, p. 809], $|s(O)|=\sum_{m=1}^{M}|s_{m}(O)|=\sum_{m=1}^{M}\rho_{m}^{n}|O|=|O|$, so that $|s(O)|=|O|$. As claimed in [19, p. 809], this implies that $s(\overline{O})=\overline{O}$. To see this, note that $s(O)\subset O$, so that $s(\overline{O})=\overline{s(O)}\subset\overline{O}$. Thus, if $s(\overline{O})\neq\overline{O}$, there exists $x\in\overline{O}\setminus s(\overline{O})$ so that, for some $y\in O$ near $x$ and some $\epsilon>0$, $B_{\epsilon}(y)\subset O\setminus s(\overline{O})\subset O\setminus s(O)$, which contradicts $|O|=|s(O)|$. Further, since $\Gamma$ is the unique fixed point of $s$, $s(\overline{O})=\overline{O}$ implies $\Gamma=\overline{O}$, and that $\Gamma$ is not disjoint then follows from Lemma 2.5. ∎
@@ 2.4 Function spaces on subsets of $\mathbb{R}^{n}$
Here we collect some results on function spaces from [17, 9, 36]. Our function spaces will be complex-valued, and we shall repeatedly use the following terminology relating to dual spaces [^4] For notational convenience we follow the convention of [17, 19] and work throughout with dual spaces of antilinear functionals rather than linear functionals.. If $X$ and $Y$ are Hilbert spaces, $X^{*}$ is the dual space of $X$, and $I:Y\to X^{*}$ is a unitary isomorphism, we say that $(Y,I)$ is a unitary realisation of $X^{*}$ and define the duality pairing $\langle\cdot,\cdot\rangle_{Y\times X}$ by $\langle y,x\rangle_{Y\times X}:=Iy(x)$, for $y\in Y,x\in X$. Having selected a unitary realisation $(Y,I)$ of $X^{*}$, we adopt $(X,I^{*})$ as our unitary realisation of $Y^{*}$, where $I^{*}x(y):=\overline{\langle y,x\rangle_{Y\times X}}$, so that $\langle x,y\rangle_{X\times Y}=\overline{\langle y,x\rangle_{Y\times X}}$, for $y\in Y,x\in X$.
@@
Similarly, $\mathbb{L}_{\infty}(\Gamma)$ denotes the space of functions on $\mathbb{R}^{n}$ that are measurable and essentially bounded with respect to $\mathcal{H}^{d}|_{\Gamma}$, normed by $\|f\|_{\mathbb{L}_{\infty}}:={\rm ess}\,\sup_{x\in\mathbb{R}^{n}}|f(x)|$. In practice we shall view $\mathbb{L}_{2}(\Gamma)$ and $\mathbb{L}_{\infty}(\Gamma)$ as spaces of functions on $\Gamma$, by identifying elements of $\mathbb{L}_{2}(\Gamma)$ and $\mathbb{L}_{\infty}(\Gamma)$ with their restrictions to $\Gamma$. The dual space $(\mathbb{L}_{2}(\Gamma))^{*}$ can be realised in the standard way as $(\mathbb{L}_{2}(\Gamma),\mathbb{I})$, where $\mathbb{I}:\mathbb{L}_{2}(\Gamma)\to\mathbb{L}_{2}(\Gamma)^{*}$ is the Riesz map (a unitary isomorphism) defined by $\mathbb{I}f(\tilde{f})=(f,\tilde{f})_{\mathbb{L}_{2}(\Gamma)}=\int_{\Gamma}f(x)\overline{\tilde{f}(x)}\,\mathrm{d}\mathcal{H}^{d}(x)$.
Now assume that $\Gamma$ is a $d$-set and that $0<d\leq n$. Then function spaces on $\mathbb{R}^{n}$ and $\Gamma$ are related via the trace operator $\mathrm{tr}_{\Gamma}$ of [50, §18.5]. Defining $\mathrm{tr}_{\Gamma}(\varphi)=\varphi|_{\Gamma}\in\mathbb{L}_{2}(\Gamma)$ for $\varphi\in C_{0}^{\infty}(\mathbb{R}^{n})$, one can show [50, Thm 18.6] that if
@@
with dense range. This trivially holds also for $d=n$, since the embedding of $H^{s}(\mathbb{R}^{n})$ into $L_{2}(\mathbb{R}^{n})$, for $s>0$, and the trace $\mathrm{tr}_{\Gamma}:L_{2}(\mathbb{R}^{n})\to\mathbb{L}_{2}(\Gamma)$ are both continuous with dense range. Setting
@@
for $t^{\prime}>t>0$, and the embeddings are continuous with dense range. As explained in [9, Rem. 6.4] ($\mathbb{H}^{t}(\Gamma)$ is denoted $\mathbb{H}^{t}_{2,0}(\Gamma)$ in [9, §6]), for the case $0<d<n$, under the further assumption that $t<1$, $\mathbb{H}^{t}(\Gamma)$ coincides (with equivalent norms) with the Besov space $B^{t}_{2,2}(\Gamma)$ of [37]. Arguing in the same way, using [37, Theorem VI.1] and that $H^{s}(\mathbb{R}^{n})$ coincides with the Besov space $B^{s}_{2,2}(\mathbb{R}^{n})$ (e.g., [37, p. 8]) [^6] Our standard notation for Besov spaces on $\mathbb{R}^{n}$ is that of, e.g., [51]., this holds also for $d=n$.
@@
2.5 Function spaces on planar screens